QUESTION 4 Quinlan Enterprises stock trades for $52.50 per share. It is expected to pay a $2.50 dividend at year end (D 1 = $2.50), and the dividend is expected to grow at a constant rate of 5.50% a year. The before-tax cost of debt is 7.50%, and the tax rate is 25%. The target capital structure consists of 45% debt and 55% common equity. What is the company's WACC if all the equity used is from reinvested earnings?

a. 7.85% b. 8.50% c. 8.84% d. 7.53% e. 8.18%

QUESTION 5

To estimate the company's WACC, Marshall Inc. recently hired you as a consultant. You have obtained the following information. (1) The firm's noncallable bonds mature in 20 years, have an 8.00% annual coupon, a par value of $1,000, and a market price of $1,050.00. (2) The company's tax rate is 25%. (3) The risk-free rate is 4.50%, the market risk premium is 5.50%, and the stock's beta is 1.20. (4) The target capital structure consists of 35% debt and the balance is common equity. The firm uses the CAPM to estimate the cost of common stock, and it does not expect to issue any new shares. What is its WACC?

a.

9.19%

b.

8.73%

c.

8.29%

d.

7.88%

e.

7.48%

QUESTION 6

Trahern Baking Co. common stock sells for $32.50 per share. It expects to earn $3.50 per share during the current year, its expected dividend payout ratio is 65%, and its expected constant dividend growth rate is 6.0%. New stock can be sold to the public at the current price, but a flotation cost of 5% would be incurred. What would be the cost of equity from new common stock?

a.

13.37%

b.

12.70%

c.

14.04%

d.

14.74%

e.

15.48%

QUESTION 7

Westbrook's Painting Co. plans to issue a $1,000 par value, 20-year noncallable bond with a 7.00% annual coupon, paid semiannually. The company's marginal tax rate is 25%, but Congress is considering a change in the corporate tax rate to 15%. By how much would the component cost of debt used to calculate the WACC change if the new tax rate was adopted?

a.

0.77%

b.

0.57%

c.

0.70%

d.

0.85%

e.

0.63%

QUESTION 8

Perpetual preferred stock from Franklin Inc. sells for $97.50 per share, and it pays an $8.50 annual dividend. If the company were to sell a new preferred issue, it would incur a flotation cost of 4.00% of the price paid by investors. What is the company's cost of preferred stock for use in calculating the WACC?

a.

9.44%

b.

8.72%

c.

9.08%

d.

10.22%

e.

9.82%
Answer #1
Weight of equity = 1-D/A
Weight of equity = 1-0.45
W(E)=0.55
Weight of debt = D/A
Weight of debt = 0.45
W(D)=0.45
Cost of equity
As per DDM
Price= Dividend in 1 year/(cost of equity - growth rate)
52.5 = 2.5/ (Cost of equity - 0.055)
Cost of equity% = 10.26
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7.5*(1-0.25)
= 5.625
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.63*0.45+10.26*0.55
WACC =8.18%
Please ask remaining parts seperately, questions are unrelated